a sphere consists of inner portion of readius r made of a material of density 8gm/cm^3 and an outer layer of thickness t made of a material of density 4 gm/cm^3. this sphere floats in a liquid of density 10gm/cm^3 with half its volume submerged .the ratio r/t is equal to ....???????????????????            plzzzzzz       help...........

Grade:Upto college level

1 Answers

Amitabh Shrivastava
13 years ago

Since half on sphere is submerged,Density(overall)of sphere=Density of liq./2.

Now density of sphere=[4/3(pi)r^3*8+4/3(pi)*((r+t)^3-r^3)*4]/4/3*(pi)*(r+t)^3 = 10/2=5

Just solve the above equation to get r/t